OpenStudy (anonymous):

Can someone help me understand this? Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0

OpenStudy (anonymous):

do u understand the interval notation [ )

OpenStudy (anonymous):

u would think where is sinx = 0 cosx = 0 at around 2pi

OpenStudy (anonymous):

No, I don't. The way my online class explains it is really hard to grasp. I was hoping someone knew a simpler way.

OpenStudy (anonymous):

i think one of them would be pi/2 cuz when sin(pi/2)*cos(pi/2) = 0 holds true

OpenStudy (anonymous):

basically 90 180 270 degrees will hold true

OpenStudy (anonymous):

which is pi/2 , pi, 3pi/2

OpenStudy (campbell_st):

well there are lots of ways to view this where does sin(x)= 0 and where does cos(x) = 0 if either factor is zero, then the product will be zero... and alternative is to use substitution sin(2x) = 2sin(x)cos(x) so you can say 1/2sin(2x) = 0 hope it helps

OpenStudy (anonymous):

u don't want to go to 2pi or more cuz that would include outside ur interval notation

OpenStudy (anonymous):

0 is also an answer

OpenStudy (anonymous):

How did you do that substitution? Where did you pull the 2x from?

OpenStudy (campbell_st):

well there is a thing called double angle trig identities one of which is sin(2x) = 2sin(x)cos(x) if you haven't covered it... just look at where cos(x) = 0 and sin(x) = 0

OpenStudy (anonymous):

They both =0 at different times.

OpenStudy (campbell_st):

thats correct... because you are multiplying sin(x)cos(x) = 0 if sin(x) = 0...... then the product of the 2 terms is zero and cos(x) = 0 ..... then the product of the 2 terms is zero thats why there well be 4 solutions.....you're interval doesn't include 2pi... which would have been a 5th solution.

OpenStudy (anonymous):

So they could be.. pi, 3pi/2, and pi/2... Would zero count?

OpenStudy (campbell_st):

yes... since the interval says state at 0 and go up to 2pi... but don't include it

OpenStudy (anonymous):

Oh, okay. So my answer will only have those three?

OpenStudy (anonymous):

Uhmm.. None of my choices have three..

OpenStudy (anonymous):

when u have [ ] it includes those in the interval when u have ( ) these are open and don't include in the interval so u have [ )

OpenStudy (campbell_st):

thats because there are 4....

OpenStudy (anonymous):

u have 0 , pi/2, pi, and 3pi/2

OpenStudy (anonymous):

Oh! SO because the zero has a [, it's included, but not 2pi, because it has a )?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

Is this one the same way? Find all solutions in the interval [0, 2π). 2 sin^2x = sin x

OpenStudy (campbell_st):

well for that subtract sin(x) from both sides of the equation then factor sin(x) sin(x)(2sin(x) - 1) = 0 now you solve each factor to find the solutions.

OpenStudy (anonymous):

How did he get that from subtracting sin x?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

took out a sinx as above was written 2sinx^2 - sinx =0 so sinx(2sinx - 1) = 0

OpenStudy (anonymous):

What do I do next? Graph them, and see where they equal zero? @fiberdust

OpenStudy (campbell_st):

if you have 2sin^2(x) = sin(x) -sin(x) -sin(x) ---------------------- 2sin^2(x) - sin(x) = 0 now factor sin(x)(2sin(x) - 1) = 0 so you need to find by solving where sin(x) = 0 and 2sin(x) - 1 = 0 hope it makes sense.